nolegirl24:

Identify the vertex for the graph of y = −3x^2 + 6x − 4. (1, 5) (1, −1) (−1, −13) (−1, −9)

nolegirl24:

@dude

dude:

To find the vertex we use the equation \(x=\large \frac{-b}{2a}\)

nolegirl24:

I thought so!! but i got x= -0.5 after i solved

dude:

Not quite, check your math \(\large\frac{-(6)}{2\times (-3)}\)

nolegirl24:

Oh i accidentally got the numbers mixed up

dude:

Ah its okay haha

nolegirl24:

x = 1 right?

dude:

Right, so we can eliminate C and D

dude:

To find the y-value we substitute 1 back into the equation \(y = −3(1)^2 + 6(1) − 4\)

nolegirl24:

okay so i got y = 3^2 + 6 - 4 y = 9 + 6 - 4 y = 15 - 4 y = 11

dude:

You made a common mistake Well I know anything times 1 is itself but you are supposed to follow PEMDAS \(y = −3(1)^2 + 6(1) − 4\) You are supposed to do \(1^2\) first giving you \(y = −3 + 6 − 4\) (Also watch your signs, -3)

nolegirl24:

y = -1?

dude:

Right

dude:

\((1,-1)\) is your answer

nolegirl24:

Okay so i had it before i just mixed up the A and B.

dude:

Yes